# Sequence Compression Speeds Up Credit Assignment in Reinforcement Learning

### v1

@@ Abstract
Temporal credit assignment in reinforcement learning is challenging due to delayed and stochastic outcomes. Monte Carlo targets can bridge long delays between action and consequence but lead to high-variance targets due to stochasticity. Temporal difference (TD) learning uses bootstrapping to overcome variance but introduces a bias that can only be corrected through many iterations. TD($\lambda$) provides a mechanism to navigate this bias-variance tradeoff smoothly. Appropriately selecting $\lambda$ can significantly improve performance. Here, we propose Chunked-TD, which uses predicted probabilities of transitions from a model for computing $\lambda$-return targets. Unlike other model-based solutions to credit assignment, Chunked-TD is less vulnerable to model inaccuracies. Our approach is motivated by the principle of history compression and ‘chunks’ trajectories for conventional TD learning. Chunking with learned world models compresses near-deterministic regions of the environment-policy interaction to speed up credit assignment while still bootstrapping when necessary. We propose algorithms that can be implemented online and show that they solve some problems much faster than conventional TD($\lambda$).  [^4] Code is available at https://github.com/Aditya-Ramesh-10/chunktd.
@@ 1 Introduction
Our experiments show that Chunked-TD can assign credit much faster than conventional TD($\lambda$) in the considered Markov decision processes with tabular value functions. We also present a variant of Chunked-TD that can use factored rewards to solve a typical hard credit assignment problem.
@@ 2.1 RL notation and definitions
Updates with $\lambda$-returns can be done in an ‘offline’ manner upon completion of episodes as computing the $\lambda$-return according to Equation 1 requires knowing the entire sequence of future states and rewards. However, it is possible to approximate updates from offline $\lambda$-returns in an online, incremental way through TD($\lambda$) which uses eligibility traces (Sutton, 1984).
@@ 3 Method
First, in Section 3.1, we present an approach that uses $n$-step targets for learning value functions based on a dynamic (and stochastic) choice of $n$ obtained through sequence compression.
@@ 3.1 $n$-step chunked targets
Figure 2: Sample-based chunked-TD for state value estimation. All next states are deterministic apart from the outcome at $x_{k}$, where $P_{\pi}(x_{k+1}|x_{k})$ is low. The states in red are a sample from our algorithm, i.e., chunking based on model probabilities that provide targets for bootstrapping. The chunked target for $s_{2}$ ($G_{2}^{C}$) is shown as an example.
@@ 3.2 Chunked-TD($\lambda$)
An interesting alternative is to use a backward model for chunking or choosing $\lambda_{t+1}$. There one would have to model $\hat{P}^{\pi}(X_{t},A_{t}|X_{t+1},A_{t+1})$ or $\hat{P}^{\pi}(X_{t}|X_{t+1})$. Intuitively, this will bootstrap more from states that may be reached by many different paths, which in turn means their value is likely to be updated many times between each visit to a particular transition. Conversely, if a state is only ever visited from a single predecessor, bootstrapping is unlikely to be helpful as the value of that state and its predecessor will always be updated together even with MC updates. In Appendix C, we show that, with careful analysis, a previously proposed algorithm by Sutton & Singh (1994) can be interpreted as using a $\lambda$-return by chunking with a backward model.  [^1] While the trace-based TDC algorithm in Sutton & Singh (1994) uses a forward model to cut traces, with adjustments to the learning rate the corresponding $\lambda$-return (or forward view) amounts to using a backward model. Please see Appendix C for further details. A crucial disadvantage of backward models is that they are policy-dependent (Chelu et al., 2020), making them harder to learn, especially in control, i.e., learning the optimal policy.
@@ 4.1 Chain-and-Split
Chain-and-Split (Figure 3) is an environment that requires adaptive values of $\lambda$ for efficient policy evaluation. It is an episodic MDP in which only the first action influences transitions and rewards. One action leads to a ‘delayed’ deterministic reward, whereas the other actions lead to stochastic consequences.
@@
The agent has $n$ actions $\{a_{1},a_{2},\dots a_{n}\}$ available at the start/root state. After the first step, there is only one action available in every state. Taking $a_{1}$ at the starting state leads to a linear chain of length $H$ which ultimately provides a deterministic reward of $0.01$. This is the optimal choice. All remaining actions lead to a common ‘parent’ state on the left ($s_{L}$), which has a zero value but is followed by noisy transitions/rewards. The state $s_{L}$ branches to one of $w$ states with equal probability. Each of the $w$ states provides a deterministic reward between -1 and 1, with the average over all states being equal to zero. We set $H=20$, $n=10$, and $w=101$ for our experiment.
@@ Results
Figure 4: The action-value gap ($\Delta Q$) between optimal action $a_{1}$ and the maximum action value among the remaining actions. The true value of $\Delta Q=0.01$. Shading indicates 95% bootstrapped confidence intervals over 10 independent trials.
@@
In contrast, MC/SARSA(1) fails due to the fact that it is not bootstrapping from the value of $s_{L}$. This results in learning entirely independent action values $\hat{Q}(start,a_{i}),i\in\{2,3\dots 10\}$, without relying on the fact that all these actions actually lead to the same next state ($s_{L}$ in Figure 3). This typically results in negative values of $\Delta Q$, as the value of some action $\hat{Q}(start,a_{i}),i\in\{2,3\dots 10\}$ is overestimated due to limited samples.
@@
Table 1 presents results for the best choice of learning rate for each approach. Results with all considered learning rates are presented in Figure 8. Expected-SARSA($\lambda$) does not reliably learn to collect the treasure. Similarly, the default version of Chunked Expected-SARSA (C-default) performs poorly.
Since rewards are suitably decomposed, we can use the factored version of Chunked Expected-SARSA (C-factored). C-factored bootstraps separately based on the predictability of each component of the state. For e.g., notice that the key, door, treasure, and time steps are perfectly predictable from the previous state. C-factored behaves like an MC update with respect to the value of only these factors. This allows the agent to better estimate Q-values quickly and successfully collect the treasure in most episodes.
@@ 5 Related Work
Unlike our model-based approach to select $\lambda$, other works have studied model-free RL algorithms with adaptive values of $\lambda$. White & White (2016) discuss several factors that could guide the choice of $\lambda$ in TD algorithms: bias-variance control, confidence in value estimates, etc. They design an objective that optimizes the immediate bias-variance trade-off to generate targets as a mixture of TD(0) and MC. Riquelme et al. (2019) propose an approach to policy evaluation that switches between TD(0) and MC based on confidence estimates of the value at that state. Watkins (1989) proposes the Q($\lambda$) algorithm, which cuts the trace whenever an exploratory action is taken. Xu et al. (2018) use meta-gradients to adjust $\lambda$.
@@
Alternatives to standard forward dynamics models can be used for assigning credit with real experience. The hindsight credit assignment (HCA) family of approaches learns a temporally extended inverse dynamics model to ascertain the influence of actions (Harutyunyan et al., 2019; Meulemans et al., 2023). Similar to backward models, temporally extended inverse dynamics models are policy-dependent and can be challenging to learn.
@@ 6 Limitations
A crucial limitation of our approach is that learning a model of the world can be challenging and expensive. Further, our approach requires a ‘generative’ model that can estimate the probability of the next state, given the current state and action. Nevertheless, our approach could be particularly useful when pre-trained generative models are already available or are being learned for imagination-based planning.
@@ 7 Conclusion
The clearest direction for future work is to extend this idea to higher dimensional state spaces and continuous actions. Other exciting next steps are to generalize our approach to the off-policy setting and explore ideas to chunk trajectories in a model-free manner.
@@ Acknowledgements
Societal Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1.2 Implementation details
We use Algorithm 2. To model $\hat{P}^{\pi}(x_{t+1},a_{t+1}|x_{t},a_{t})$ in this case, we simply maintain empirical counts of each possible transition $n(s,a,s^{\prime})$ and state-action pair $n(s,a)$ and let $\hat{P}(x_{t+1}|x_{t},a_{t})=\frac{n(s_{t},a_{t},s_{t+1})}{n(s_{t},a_{t})}$. We decompose $\hat{P}^{\pi}(x_{t+1},a_{t+1}|x_{t},a_{t})=\hat{P}(x_{t+1}|x_{t},a_{t})\pi(a_{t+1}|s_{t+1})$
@@ A.2.1 Environment details
During the episode, at $k$ values of the time step, the agent stochastically accumulates ‘charge’. The $k$ time steps where a charge is accumulated are fixed for the interaction but are randomly chosen for each independent initialization (seed) of the environment. In our selected setting, the episode lasts $200$ steps, and the agent encounters $k=10$ accumulation points, where it accumulates charge sampled from a binomial variable with $n=H/k=20$ and $p=0.5$.
@@ A.2.2 Implementation details
Since the true state space is quite large, with 200 possible values of charge and time step, we learn a neural network model to predict $\Delta s_{t+1}=s_{t+1}-s_{t}$ as a categorical distribution. Using state-delta has the benefit of reducing the space of outputs under the model and is similar to models learned in video prediction. The probability of the whole next state (or state-delta) is taken as the product of the probability of each individual component (each component is a categorical variable in $n=H/k$, the maximum change possible in a state-component in this environment). We use the probability of the actual delta and use $\hat{P}(\Delta s_{t+1}|s_{t},a_{t})$ as the model’s predicted probability of the transition $\lambda_{t+1}=\hat{P}(\Delta s_{t+2}|s_{t+1},a_{t+1})\pi(a_{t+1}|s_{t+1})$.
@@ A.3.2 Implementation details
The C-factored algorithm refers to Algorithm 4. We decompose $\hat{P}^{\pi}(x^{i}_{t+1},|x_{t})=\sum_{a}\hat{P}(x^{i}_{t+1}|x_{t},a)\pi(a|s_{t+1}).$ Recall that $x_{t}=s_{t}$ in our experiments.
@@ B.2 Chunked SARSA and Chunked Expected-SARSA
An advantage of the Chunked Expected-SARSA over Chunked SARSA is in situations where multiple actions lead to the same next state. Notice that Chunked SARSA would cut the trace when an action of low probability is taken, even if the next state is the same for all actions ($\lambda_{t}=\hat{P}(X_{t+1}|X_{t},A_{t})\pi(A_{t+1}|X_{t+1})$). This would not happen with Chunked Expected-SARSA, as it averages over the consequence of all actions, i.e., $\lambda_{t}=\sum_{a}\hat{P}(X_{t+1}|X_{t},a)\pi(a|X_{t+1})$.
@@ B.3 Component-wise chunked-TD
7:     $\delta_{t}\leftarrow R_{t+1}+\hat{V}(S_{t+1})-\hat{Q}(S_{t},A_{t})$, where $\hat{V}(S_{t+1})=\sum_{a}\pi(a|S_{t+1})\hat{Q}(S_{t+1},a)$
@@
8:        $\delta^{i}_{t}\leftarrow R^{i}_{t+1}+\hat{V}^{i}(S_{t+1})-\hat{Q}^{i}(S_{t},A_{t})$, where $\hat{V}^{i}(S_{t+1})=\sum_{a}\pi(a|S_{t+1})\hat{Q}^{i}(S_{t+1},a)$
@@ Appendix C Algorithms from Sutton and Singh 1994
where $\hat{V}_{t^{\prime}}(s)$ represents the value estimate for $s$ after accounting for the transition observed at $t^{\prime}$, $n_{t^{\prime}}(s)$ is the total number of times $s$ has been visited, and $\tau(t,t^{\prime})$ represents some time future time compared to $t$ to be defined. The approaches differ only in $\tau(t,t^{\prime})$, that is which time step’s value estimator is used to evaluate the successor states. Equation 7 is only enforced at the end of episodes for those states $s$ which occur within each episode. The values of states which do not occur within an episode are left unchanged. Note that here, we use a shared time index across all episodes such that if an episode ends at time $t$, the next episode will begin at $t+1$. We will assume acyclic MDPs as in Sutton & Singh (1994).
@@ C.1 Naive TD(0)
which just computes an incremental average of the value estimates and rewards observed directly following each state $S$. This is easily seen to be $TD(0)$ with a learning rate of $\frac{1}{n(S_{t})}$ and we can identify the $\lambda$ return as $G_{t}^{\lambda}=R_{t+1}+\hat{V}(S_{t+1})$. This, however, suffers from only backing up reward information one step at a time.
@@ C.2 TD(1/n)
$\tilde{V}_{t}$ $\leftarrow\hat{V}(S_{t})+\frac{1}{n(S_{t})}(R_{t+1}+\tilde{V}_{t+1}-\hat{V}(S_{t})),$ (11)
@@
$\implies G_{t}^{\lambda}$ $=R_{t+1}+\frac{1}{n(S_{t+1})}G_{t+1}^{\lambda}+\frac{n(S_{t+1})-1}{n(S_{t+1})}\hat{V}(S_{t+1})).$
@@
We can confirm that the TD$(1/n)$ algorithm (presented as Algorithm 5) achieves the update of Equation 11 in an incremental manner.
@@
3:     $e({S_{t}})\leftarrow 1$increment trace for current state}
@@
7:        $e(s)\leftarrow e(s)\frac{1}{n(S_{t})}$decay traces}

### v2

@@ Abstract
Temporal credit assignment in reinforcement learning is challenging due to delayed and stochastic outcomes. Monte Carlo targets can bridge long delays between action and consequence but lead to high-variance targets due to stochasticity. Temporal difference (TD) learning uses bootstrapping to overcome variance but introduces a bias that can only be corrected through many iterations. TD($\lambda$) provides a mechanism to navigate this bias-variance tradeoff smoothly. Appropriately selecting $\lambda$ can significantly improve performance. Here, we propose Chunked-TD, which uses predicted probabilities of transitions from a model for computing $\lambda$-return targets. Unlike other model-based solutions to credit assignment, Chunked-TD is less vulnerable to model inaccuracies. Our approach is motivated by the principle of history compression and ‘chunks’ trajectories for conventional TD learning. Chunking with learned world models compresses near-deterministic regions of the environment-policy interaction to speed up credit assignment while still bootstrapping when necessary. We propose algorithms that can be implemented online and show that they solve some problems much faster than conventional TD($\lambda$).
@@ 1 Introduction
Our experiments show that Chunked-TD can assign credit much faster than conventional TD($\lambda$) in the considered Markov decision processes with tabular value functions. We also present a variant of Chunked-TD that can use factored rewards to solve a typical hard credit assignment problem.  [^4] Code is available at https://github.com/Aditya-Ramesh-10/chunktd.
@@ 2.1 RL notation and definitions
Updates with $\lambda$-returns can be done in an offline manner upon completion of episodes. This is because computing the $\lambda$-return according to Equation 1 requires knowing the entire sequence of future states and rewards. However, it is possible to approximate updates from offline $\lambda$-returns in an online, incremental way through TD($\lambda$), which uses eligibility traces (Sutton, 1984).
@@ 3 Method
First, in Section 3.1, we present an approach that uses $n$-step targets for learning value functions based on a dynamic (and stochastic) choice of $n$ obtained through sequence compression/chunking.
@@ 3.1 $n$-step chunked targets
Figure 2: Sample-based chunking (see Section 3.1) for state value estimation. All next states (and rewards) are deterministic apart from the outcome at $x_{k}$, where $P_{\pi}(x_{k+1}|x_{k})$ is low. The states in red are a sample from our algorithm, i.e., chunking based on model probabilities that provide targets for bootstrapping. The chunked target for $s_{2}$ (with $\gamma=1$, denoted as $G_{2}^{C}$) is shown as an example.
@@ 3.2 Chunked-TD($\lambda$)
An interesting alternative is to use a backward model for chunking or choosing $\lambda_{t+1}$. There one would have to model $\hat{P}^{\pi}(X_{t},A_{t}|X_{t+1},A_{t+1})$ or $\hat{P}^{\pi}(X_{t}|X_{t+1})$. Intuitively, this will bootstrap more from states that may be reached by many different paths, which in turn means their value is likely to be updated many times between each visit to a particular transition. Conversely, if a state is only ever visited from a single predecessor, bootstrapping is unlikely to be helpful as the value of that state and its predecessor will always be updated together, even with MC updates. In Appendix C, we show that, with careful analysis, a previously proposed algorithm by Sutton & Singh (1994) can be interpreted as using a $\lambda$-return by chunking with a backward model.  [^2] While the trace-based TDC algorithm in Sutton & Singh (1994) uses a forward model to cut traces, with adjustments to the learning rate the corresponding $\lambda$-return (or forward view) amounts to using a backward model. Please see Appendix C for further details. A crucial disadvantage of backward models is that they are policy-dependent (Chelu et al., 2020), making them harder to learn, especially in control, i.e., learning the optimal policy.
@@ 4.1 Chain-and-Split
Chain-and-Split (see Figure 3) is an environment that requires adaptive values of $\lambda$ for efficient policy evaluation. It is an episodic MDP in which only the first action influences transitions and rewards. One action leads to a ‘delayed’ deterministic reward, whereas the other actions lead to stochastic consequences.
@@
The agent has $n$ actions $\{a_{1},a_{2},\dots a_{n}\}$ available at the start/root state. After the first step, there is only one action available in every state. Taking $a_{1}$ at the start state leads to a linear chain of length $H$, which ultimately provides a deterministic reward of $0.01$. This is the optimal choice. All remaining actions lead to a common ‘parent’ state on the left ($s_{L}$), which has a zero value but is followed by noisy transitions/rewards. The state $s_{L}$ branches to one of $w$ states with equal probability. Each of the $w$ states provides a deterministic reward between -1 and 1, with the average over all states being equal to zero. We set $H=20$, $n=10$, and $w=101$ for our experiment.
@@ Results
Figure 4: Results from the Chain-and-Split environment. The action-value gap ($\Delta Q$) between optimal action $a_{1}$ and the maximum action value among the remaining actions. The true value of $\Delta Q=0.01$. Shading indicates 95% bootstrapped confidence intervals over 10 independent trials.
@@
In contrast, MC/SARSA(1) fails because it does not bootstrap from the value of $s_{L}$. This results in learning entirely independent action values $\hat{Q}(start,a_{i}),i\in\{2,3\dots 10\}$, without relying on the fact that all these actions actually lead to the same next state ($s_{L}$ in Figure 3). This results in negative values of $\Delta Q$, as the value of some action $\hat{Q}(start,a_{i}),i\in\{2,3\dots 10\}$ is typically overestimated due to limited samples.
@@
Table 1 presents results for the best choice of learning rate for each approach. Results with all considered learning rates are presented in Figure 8. Expected-SARSA($\lambda$), with $\lambda\in\{0,0.1,0.5,0.9,1\}$, does not reliably learn to collect the treasure. Similarly, the default version of Chunked Expected-SARSA (C-default) performs poorly.
Since rewards are suitably decomposed, we could use the factored version of Chunked Expected-SARSA (C-factored). C-factored bootstraps separately based on the predictability of each component of the state. For e.g., notice that the key, door, treasure, and time steps are perfectly predictable from the previous state. C-factored behaves like an MC update with respect to the value of only these factors. This allows the agent to better estimate Q-values quickly and successfully collect the treasure in most episodes.
@@ 5 Related Work
Unlike our model-based approach to select $\lambda$, other works have studied model-free RL algorithms with adaptive or state-dependent values of $\lambda$. White & White (2016) discuss several factors that could guide the choice of $\lambda$ in TD algorithms: bias-variance control, confidence in value estimates, etc. They design an objective that optimizes the immediate bias-variance trade-off to generate targets as a mixture of TD(0) and MC. Riquelme et al. (2019) propose an approach to policy evaluation that switches between TD(0) and MC based on confidence estimates of the value at that state. Watkins (1989) proposes the Q($\lambda$) algorithm, which cuts the trace whenever an exploratory action is taken. Xu et al. (2018) use meta-gradients to adjust $\lambda$.
@@
Alternatives to standard forward dynamics models can be used for assigning credit with real experience. The hindsight credit assignment (HCA) family of approaches learns a temporally extended inverse dynamics model to ascertain the influence of actions (Harutyunyan et al., 2019; Alipov et al., 2021; Meulemans et al., 2023). Similar to backward models, temporally extended inverse dynamics models are policy-dependent and can be challenging to learn.
@@ 6 Limitations
A crucial limitation of our approach is that learning a model of the world can be challenging and expensive. Furthermore, our approach requires a ‘generative’ model that can estimate the probability of the next state, given the current state and action. Nevertheless, our approach could be particularly useful when pre-trained generative models are already available or are being learned for imagination-based planning.
@@ 7 Conclusion
The clearest direction for future work is to extend this idea to higher dimensional state spaces and continuous actions. Other exciting next steps are to generalize our approach to the off-policy setting and to further explore different methods of chunking trajectories, e.g., with policy-conditioned backward models or even in model-free ways.
@@ Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1.2 Implementation details
We use Algorithm 2. To model $\hat{P}^{\pi}(x_{t+1},a_{t+1}|x_{t},a_{t})$ in this case, we simply maintain empirical counts of each possible transition $n(x,a,x^{\prime})$ and state-action pair $n(x,a)$ and let $\hat{P}(x_{t+1}|x_{t},a_{t})=\frac{n(x_{t},a_{t},x_{t+1})}{n(x_{t},a_{t})}$. We decompose $\hat{P}^{\pi}(x_{t+1},a_{t+1}|x_{t},a_{t})=\hat{P}(x_{t+1}|x_{t},a_{t})\pi(a_{t+1}|s_{t+1})$. Recall that $x_{t}$ is the same as $s_{t}$ in our experiments.
@@ A.2.1 Environment details
During the episode, at $k$ values of the time step, the agent stochastically accumulates ‘charge’. The $k$ time steps where a stochastic charge is accumulated are fixed for the interaction (all episodes) but are randomly chosen for each independent initialization (seed) of the environment. In our selected setting, the episode lasts $200$ steps, and the agent encounters $k=10$ accumulation points, where it accumulates charge sampled from a binomial variable with $n=H/k=20$ and $p=0.5$.
@@ A.2.2 Implementation details
Since the state space is quite large, with 200 possible values of charge and time step, we learn a neural network model to predict $\Delta s_{t+1}=s_{t+1}-s_{t}$ as a categorical distribution. Using state-delta has the benefit of reducing the space of outputs under the model and is similar to models learned in video prediction. The probability of the whole next state (or state-delta) is modeled independently, i.e., taken as the product of the probability of each individual component (each component is a categorical variable in $n=H/k$, the maximum change possible in a state component in this environment). We use the probability of the actual delta and use $\hat{P}(\Delta s_{t+1}|s_{t},a_{t})$ as the model’s predicted probability of the transition $\lambda_{t+1}=\hat{P}(\Delta s_{t+2}|s_{t+1},a_{t+1})\pi(a_{t+1}|s_{t+1})$.
@@ A.3.2 Implementation details
The C-factored algorithm refers to Algorithm 4. We decompose $\hat{P}^{\pi}(x^{i}_{t+1},|x_{t})=\sum_{a}\hat{P}(x^{i}_{t+1}|x_{t},a)\pi(a|s_{t}).$ Recall that $x_{t}=s_{t}$ in our experiments.
@@ B.2 Chunked SARSA and Chunked Expected-SARSA
An advantage of the Chunked Expected-SARSA over Chunked SARSA is in situations where multiple actions lead to the same next state. Notice that Chunked SARSA would cut the trace when an action of low probability is taken, even if the next state is the same for all actions ($\lambda_{t}=\hat{P}(X_{t+1}|X_{t},A_{t})\pi(A_{t+1}|X_{t+1})$). This would not happen with Chunked Expected-SARSA, as it averages over the consequence of all actions, i.e., $\lambda_{t}=\sum_{a}\hat{P}(X_{t+1}|X_{t},a)\pi(a|X_{t})$.
@@ B.3 Component-wise chunked-TD
7:     $\delta_{t}\leftarrow R_{t+1}+\gamma\hat{V}(S_{t+1})-\hat{Q}(S_{t},A_{t})$, where $\hat{V}(S_{t+1})=\sum_{a}\pi(a|S_{t+1})\hat{Q}(S_{t+1},a)$
@@
8:        $\delta^{i}_{t}\leftarrow R^{i}_{t+1}+\gamma\hat{V}^{i}(S_{t+1})-\hat{Q}^{i}(S_{t},A_{t})$, where $\hat{V}^{i}(S_{t+1})=\sum_{a}\pi(a|S_{t+1})\hat{Q}^{i}(S_{t+1},a)$
@@ Appendix C Algorithms from Sutton and Singh 1994
where $\hat{V}_{t^{\prime}}(s)$ represents the value estimate for $s$ after accounting for the transition observed at $t^{\prime}$, $n_{t^{\prime}}(s)$ is the total number of times $s$ has been visited, and $\tau(t,t^{\prime})$ represents some time future time compared to $t$ to be defined. The approaches differ only in $\tau(t,t^{\prime})$, that is, which time step’s value estimator is used to evaluate the successor states. Equation 7 is only enforced at the end of episodes for those states $s$ which occur within each episode. The values of states which do not occur within an episode are left unchanged. Note that here, we use a shared time index across all episodes, so if an episode ends at time $t$, the next episode will begin at $t+1$. We will assume acyclic MDPs as in Sutton & Singh (1994).
@@ C.1 Naive TD(0)
which just computes an incremental average of the value estimates and rewards observed directly following each state $S$. This is easily seen to be $TD(0)$ with a learning rate of $\frac{1}{n(S_{t})}$ and we can identify the target as $G_{t}^{\lambda}=R_{t+1}+\hat{V}(S_{t+1})$. This, however, suffers from only backing up reward information one step at a time.
@@ C.2 TD(1/n)
$\tilde{V}_{t}$ $\leftarrow\hat{V}(S_{t})+\frac{1}{n(S_{t})}(R_{t+1}+\tilde{V}_{t+1}-\hat{V}(S_{t})).$ (11)
@@
$\implies G_{t}^{\lambda}$ $=R_{t+1}+\frac{1}{n(S_{t+1})}G_{t+1}^{\lambda}+\frac{n(S_{t+1})-1}{n(S_{t+1})}\hat{V}(S_{t+1}).$
@@
We can confirm that the TD$(1/n)$ algorithm (Algorithm 5) achieves the update of Equation 11 in an incremental manner.
@@
3:     $e({S_{t}})\leftarrow 1$
@@
7:        $e(s)\leftarrow e(s)\frac{1}{n(S_{t})}$